a hammer of mass 1 kg strikes on the head of a nail with velocity of 2ms-1. it drives the nail 0.01 m into a wooden block. find the force applied by the hammer and the time of impact.

Grade:11

3 Answers

Aziz Alam IIT Roorkee
9 years ago
v=0 m/s , u=2 m/s => v2-u2 =2as => a=200 m/s2

Force applied by hammer= 200x1=200 N

time of impact= 2/200=0.01 seconds=10 milli seconds


Thanks & Regards
Aziz Alam
IIT Roorkee
askIITians
saket
7 years ago
A uniform rope of length L,resting on a frictionless horizontal table is pulled at one end by a force F.What is the tension in the rope at a distance x from the end where the force is applied?
SAHIL
7 years ago
Let n be the mass per unit length of rope. Therefore, mass of rope = nL
. Acceleration in the rope due to force F will be a = F/nL.
Mass of rope of length (L - l ) will be n (L - l ).
Therefore, tension in the rope of length (L - l ), is equal to pulling force on it = n (L - l ) a = n (L - l ) × F/nL = F (1 - el /L)